# Controllable Human-centric Keyframe Interpolation with Generative Prior

### v1

@@ Abstract
Existing interpolation methods use pre-trained video diffusion priors to generate intermediate frames between sparsely sampled keyframes. In the absence of 3D geometric guidance, these methods struggle to produce plausible results for complex, articulated human motions and offer limited control over the synthesized dynamics. In this paper, we introduce PoseFuse3D Keyframe Interpolator (PoseFuse3D-KI), a novel framework that integrates 3D human guidance signals into the diffusion process for Controllable Human-centric Keyframe Interpolation (CHKI). To provide rich spatial and structural cues for interpolation, our PoseFuse3D, a 3D-informed control model, features a novel SMPL-X encoder that transforms 3D geometry and shape into the 2D latent conditioning space, alongside a fusion network that integrates these 3D cues with 2D pose embeddings. For evaluation, we build CHKI-Video, a new dataset annotated with both 2D poses and 3D SMPL-X parameters. We show that PoseFuse3D-KI consistently outperforms state-of-the-art baselines on CHKI-Video, achieving a 9% improvement in PSNR and a 38% reduction in LPIPS. Comprehensive ablations demonstrate that our PoseFuse3D model improves interpolation fidelity.
@@ 1 Introduction
Frame interpolation aims to generate new frames between two consecutive video frames to improve temporal smoothness. Traditional interpolation methods [11, 43, 16] assume small, simple motion over short time spans. These methods are challenged when the input frames are widely separated, known as keyframe interpolation or generative inbetweening [52, 44, 39, 10], where the motion between them becomes complex and ambiguous. This challenge is magnified in human-centric videos, where articulated body movements encompass diverse poses and shapes. With human subjects being prevalent in today’s video content, there is a growing need for interpolation methods to handle large temporal gaps and intricate human motion while offering plausible results.
@@
Human-centric keyframe interpolation remains challenging for current methods. Recent approaches [10, 44, 39] leverage generative priors from image-to-video (I2V) models to bridge the temporal gap. These methods condition the interpolation solely on the input keyframes without intermediate guidance. Consequently, they often struggle to resolve motion ambiguities and accurately capture the complex articulated dynamics of human motion. For instance, when keyframes involve large occlusions or non-rigid joint movements, these methods often produce implausible or distorted interpolations (Figure 1(a)) due to insufficient intermediate guidance. FCVG [52] has explored interpolation keyframes with 2D skeletons as control signals for human subjects. However, 2D lines cannot convey full body shape and geometry, leading to unrealistic results (see Figure 1(b)). These methods lack fine-grained control over the interpolation process, limiting their ability to produce flexible, high-fidelity human-centric interpolations.
@@
In this study, we investigate the integration of 3D human conditions into the human-centric keyframe interpolation pipeline. Drawing inspiration from recent advances in human animation [51, 50], we propose to integrate 2D human poses [45] with 3D SMPL-X models [28] as intermediate control signals. These signals provide precise guidance for complex articulated motions: 2D poses offer concise representations of human joint poses, while 3D models capture rich spatial geometry. However, effectively processing these control signals poses challenges. First, common practice renders 3D human models into 2D proxies (e.g., colored surface, normals, depth maps) before encoding, leading to substantial loss of spatial information in occluded regions. Therefore, we need to develop a dedicated encoder that preserves occluded 3D details when converting models into control signals. Second, fusing signals with different information content and granularity is nontrivial. This necessitates designing appropriate neural architectures that can accurately extract 3D cues and harmonize them with 2D poses into a unified, informative control input.
@@
To this end, we introduce PoseFuse3D Keyframe Interpolator, termed as PoseFuse3D-KI, a novel pose-control framework for controllable human-centric keyframe interpolation (Figure 1(c)). Our framework is unique in its 3D-informed control model PoseFuse3D, which comprises three jointly trained modules. The first derives control features from visualized conditions; the second encodes and aggregates 3D body geometry into 2D image conditioning features; and the third combines outputs from the first two modules into a unified control signal for interpolation. In contrast to Champ [51], which fuses 3D features from rendered visualizations, our encoder processes features directly in 3D and integrates projected features through feature aggregation.
To evaluate the proposed PoseFuse3D-KI, we have also built a high-quality video dataset for Controllable Human-centric Keyframe Interpolation (CHKI). Most existing interpolation datasets, such as SportsSlomo [5], target small temporal gaps, lack annotations for 2D poses or 3D human models, and offer limited human-centric motion diversity. Therefore, we introduce CHKI-Video, a new dataset for a systematic evaluation of CHKI algorithms. CHKI-Video comprises 2,614 high-quality video clips of over 180K frames sourced from SportsSlomo [5] and Pexels [^1] https://www.pexels.com/videos/ website that hosts high-quality stock videos. Each frame is carefully annotated with bounding boxes, segmentation masks, 2D human poses, and SMPL-X parameters, using state-of-the-art tools [3, 18, 23, 32, 40] supplemented by manual verification. From this collection, we derive a benchmark specifically for the CHKI task. We hope this benchmark will help improve controlled keyframe interpolation techniques with its high-quality videos and diverse examples.
Our contributions are threefold: (i) We present PoseFuse3D-KI, an effective interpolation framework for human-centric keyframe interpolation, characterized by a novel pose-control model, PoseFuse3D. It effectively extracts control signals from 3D SMPL-X and fuses 2D signals, allowing precise and informative control. (ii) For evaluation, we construct CHKI-Video, a benchmark dataset with comprehensive human-centric annotations, which are absent in existing interpolation benchmarks. (iii) Through extensive experiments, we demonstrate that PoseFuse3D-KI delivers state-of-the-art performance on our CHKI-Video benchmark with an improvement of 1.85 dB in PSNR and a reduction of 0.0796 in LPIPS.
@@ 2 Related Work
Frame Interpolation. Traditional frame interpolation methods are primarily designed for temporally adjacent frames and rely on either direct synthesis using convolutional networks [7, 17], or motion representations such as dynamic kernels [26, 27, 30, 6, 8, 20] and optical flows [15, 21, 19, 35, 13, 1, 11, 16, 41, 46]. Recent advances [10, 44, 39] extend this task to more challenging keyframe interpolation scenarios by leveraging the generative priors of image-to-video diffusion models [2, 37, 33, 12]. These methods combine temporal forward and reverse denoising predictions in a unified process to enable interpolation. However, such approaches struggle when faced with complex articulated motions or ambiguous transitions from human keyframes. To alleviate motion ambiguity in interpolation, FCVG [52] introduces 2D matched lines as control signals. However, this control signal lacks the 3D geometric context required for plausible human-centric interpolation, resulting in unrealistic body shape and appearance.
@@
Pose-Guided Human Animation. Recent advances in pose-guided human animation [50, 51, 14, 31, 47, 42, 38] harness the power of diffusion models and have achieved remarkable success in generating videos from a single reference image. These methods offer flexible and precise control by incorporating enriched conditioning signals and increasingly sophisticated control mechanisms. 2D human poses, such as OpenPose [4] and DWPose [45], are widely used in existing works [51, 50, 31, 47], but they are limited in capturing fine-grained geometry and motion dynamics. To address this, recent works [51, 50] integrate 3D human parametric models [25, 28], which offer realistic body representation through blend shapes and skinning, resulting in more accurate and expressive human animations. Specifically, Champ [51] renders 3D human models into 2D proxies (e.g., normals, depth, and semantic maps) and combines them with 2D pose visualizations as control input. It operates directly on these visualizations and unifies them using a simple summation operation. To incorporate control signals into diffusion models, many approaches [38, 42, 47, 50] adapt ControlNet [47] for their customized control networks. Some methods [51, 14] introduce task-specific pose guiders but often require retraining the majority of denoiser parameters. ControlNeXt [31] improves control efficiency by encoding conditions with a lightweight convolutional network and injecting them via cross-normalization after the first denoising block, tuning only a minimal subset of parameters. This efficient mechanism enables robust conditioning over large-scale pre-trained video generators [37, 2].
@@
In this paper, we demonstrate the advantage of combining 2D poses (DWPose [45]) with 3D human models (SMPL-X [28]) through comprehensive experiments. We present a novel pose control model that extracts unified, 3D-informed control features to provide precise guidance. Instead of relying on rendered normals and depth maps, the model directly extracts explicit 3D information from human models in 3D space, preserving richer control signals. It adopts a ControlNeXt-inspired strategy to control video diffusion models for keyframe interpolation.
@@ 3.1 PoseFuse3D
PoseFuse3D is a 3D-informed control model that provides 3D human structure and geometry guidance for plausible human interpolation. This 3D-informed guidance is injected into the base diffusion model after the first denoising block via cross-normalization [31]. Internally, PoseFuse3D comprises three jointly trained components: a visual encoding module that derives control features from 3D SMPL-X [28] renderings and 2D DWPose [45] visualizations; a SMPL-X encoder that directly embeds 3D humans and aggregates them into image conditioning maps; and a fusion module that integrates encoding streams into a unified control tensor to guide interpolation in the base VDM.
@@
Visual Encoding. This visual encoding module extracts conditional features from visualized control images, maintaining natural pixel-level alignment with the controlled video latents. To enhance the control signals, we incorporate visualizations from both DWPose and SMPL-X. While SMPL-X renderings provide rich human surface details, their keypoint information is indirect, mixed with other vertices and mesh faces. Therefore, we add DWPose visualizations to emphasize the skeletal keypoint layout for robust pose understanding. This combination was also demonstrated to be effective in prior work [51]. Specifically, our visual encoding module employs two parallel convolutional encoders to capture comprehensive pose information. One encoder processes DWPose visualizations to capture compact pose information, while the other handles SMPL-X renderings that retain 3D cues such as occlusion boundaries and projected shapes. Notably, to enrich semantic detail, we use SMPL-X Colored Surface [50] during rendering, which assigns a unique color to each vertex. The resulting feature maps, $E^{2D}$ and $E^{3D}$, are then passed to the fusion module for unified conditioning.
SMPL-X Encoder. Although 3D model renderings offer aligned image maps for conditioning in VDMs, the rendering operation discards parts of 3D information, particularly in occluded regions. This results in implausible interpolation for keyframes of large human motion. For example, in the first case of Figure 3, all variants relying solely on rendered controls fail to interpolate the occluded arms with correct spatial position. To enhance controllability with direct 3D information, we introduce a SMPL-X encoder (Figure 2(b)) that processes the SMPL-X model in 3D space and transforms it into image conditioning feature $S^{3D}$. Specifically, a SMPL-X model [28] is parameterized with $(\theta,\beta,\psi)$, corresponding to pose, shape, and expression. We obtain structural information by forwarding these parameters into the SMPL-X model to generate vertex and joint coordinates $P^{3D},J^{3D}$ in 3D space and obtain their corresponding 2D coordinates $P^{2D},J^{2D}$ through projection:
@@
where $Q\in\mathbb{R}^{B\times HW\times d}$ is the flattened $d$-dimension feature of a standard 2D meshgrid extracted by a convolutional encoder. Finally, the outputs $O^{J}$ and $O^{P}$ are concatenated and passed through a downsampling block to produce the final SMPL-X control representation $S^{3D}$, which serves as an informative and compact image embedding of the underlying 3D human structure.
Condition Fusion. The condition fusion module combines control features from both 2D and 3D signals into a unified representation to guide keyframe interpolation. For robust feature representation, we introduce a coarse-to-fine fusion strategy that progressively integrates rich geometric information from the 3D features into the compact 2D pose features. Specifically, we adopt two attention-based fusion blocks to perform this integration, where each block contains three attention layers for progressive refinement. The first layer is a self-attention module that processes the 3D features by operating on the sum of rendering encoding $E^{3D}$ and SMPL-X features $S^{3D}$. The second layer performs cross-attention, aligning the 3D features with the 2D encoding $E^{2D}$ through a spatially localized interaction scheme. Notably, we adapt the shifted window-partition strategy [24, 46] to restrict attention computation to adjacent regions, enhancing local alignment. The third layer applies temporal self-attention to capture temporal dynamic correlations in the fused representation. We use the second fusion block output as the final control signal, which is injected into the base interpolation engine to provide fine-grained, structure-aware guidance during synthesis.

### v2

@@ Abstract
Existing interpolation methods use pre‑trained video diffusion priors to generate intermediate frames between sparsely sampled keyframes. In the absence of 3D geometric guidance, these methods struggle to produce plausible results for complex, articulated human motions and offer limited control over the synthesized dynamics. In this paper, we introduce PoseFuse3D Keyframe Interpolator (PoseFuse3D-KI), a novel framework that integrates 3D human guidance signals into the diffusion process for Controllable Human-centric Keyframe Interpolation (CHKI). To provide rich spatial and structural cues for interpolation, our PoseFuse3D, a 3D‑informed control model, features a novel SMPL‑X encoder that transforms 3D geometry and shape into the 2D latent conditioning space, alongside a fusion network that integrates these 3D cues with 2D pose embeddings. For evaluation, we build CHKI-Video, a new dataset annotated with both 2D poses and 3D SMPL‑X parameters. We show that PoseFuse3D-KI consistently outperforms state-of-the-art baselines on CHKI-Video, achieving a 9% improvement in PSNR and a 38% reduction in LPIPS. Comprehensive ablations demonstrate that our PoseFuse3D model improves interpolation fidelity.
footnotetext: Corresponding authors
@@ 1 Introduction
Frame interpolation aims to generate new frames between two consecutive video frames to improve temporal smoothness. Traditional interpolation methods [guo2025generalizable, xue2019vimeo90k, jiang2018superslomo] assume small, simple motion over short time spans. These methods are challenged when the input frames are widely separated, known as keyframe interpolation or generative inbetweening [zhu2024generative, yang2024vibidsampler, wang2024generative, feng2024trf], where the motion between them becomes complex and ambiguous. This challenge is magnified in human‑centric videos, where articulated body movements encompass diverse poses and shapes. With human subjects being prevalent in today’s video content, there is a growing need for interpolation methods to handle large temporal gaps and intricate human motion while offering plausible results.
@@
Human‑centric keyframe interpolation remains challenging for current methods. Recent approaches [feng2024trf, yang2024vibidsampler, wang2024generative] leverage generative priors from image‑to‑video (I2V) models to bridge the temporal gap. These methods condition the interpolation solely on the input keyframes without intermediate guidance. Consequently, they often struggle to resolve motion ambiguities and accurately capture the complex articulated dynamics of human motion. For instance, when keyframes involve large occlusions or non‑rigid joint movements, these methods often produce implausible or distorted interpolations (Figure 1(a)) due to insufficient intermediate guidance. FCVG [zhu2024generative] has explored interpolation keyframes with 2D skeletons as control signals for human subjects. However, 2D lines cannot convey full body shape and geometry, leading to unrealistic results (see Figure 1(b)). These methods lack fine-grained control over the interpolation process, limiting their ability to produce flexible, high‑fidelity human‑centric interpolations.
@@
In this study, we investigate the integration of 3D human conditions into the human‑centric keyframe interpolation pipeline. Drawing inspiration from recent advances in human animation [zhu2024champ, zhou2024realisdance], we propose to integrate 2D human poses [yang2023dwpose] with 3D SMPL-X models [pavlakos2019smplx] as intermediate control signals. These signals provide precise guidance for complex articulated motions: 2D poses offer concise representations of human joint poses, while 3D models capture rich spatial geometry. However, effectively processing these control signals poses challenges. First, common practice renders 3D human models into 2D proxies (e.g., colored surface, normals, depth maps) before encoding, leading to substantial loss of spatial information in occluded regions. Therefore, we need to develop a dedicated encoder that preserves occluded 3D details when converting models into control signals. Second, fusing signals with different information content and granularity is nontrivial. This necessitates designing appropriate neural architectures that can accurately extract 3D cues and harmonize them with 2D poses into a unified, informative control input.
@@
To this end, we introduce PoseFuse3D Keyframe Interpolator, termed as PoseFuse3D-KI, a novel pose-control framework for controllable human-centric keyframe interpolation (Figure 1(c)). Our framework is unique in its 3D-informed control model PoseFuse3D, which comprises three jointly trained modules. The first derives control features from visualized conditions; the second encodes and aggregates 3D body geometry into 2D image conditioning features; and the third combines outputs from the first two modules into a unified control signal for interpolation. In contrast to Champ [zhu2024champ], which fuses 3D features from rendered visualizations, our encoder processes features directly in 3D and integrates projected features through feature aggregation.
To evaluate the proposed PoseFuse3D-KI, we have also built a high-quality video dataset for Controllable Human-centric Keyframe Interpolation (CHKI). Most existing interpolation datasets, such as SportsSlomo [chen2024sportsslomo], target small temporal gaps, lack annotations for 2D poses or 3D human models, and offer limited human‑centric motion diversity. Therefore, we introduce CHKI-Video, a new dataset for a systematic evaluation of CHKI algorithms. CHKI-Video comprises 2,614 high-quality video clips of over 180K frames sourced from SportsSlomo [chen2024sportsslomo] and Pexels [pexels] website that hosts high-quality stock videos. Each frame is carefully annotated with bounding boxes, segmentation masks, 2D human poses, and SMPL‑X parameters, using state‑of‑the‑art tools [cai2023smpler, khirodkar2024sapiens, liu2024groundingdino, ravi2024sam2, wu2023dover] supplemented by manual verification. From this collection, we derive a benchmark specifically for the CHKI task. We hope this benchmark will help improve controlled keyframe interpolation techniques with its high-quality videos and diverse examples.
Our contributions are threefold: (i) We present PoseFuse3D-KI, an effective interpolation framework for human-centric keyframe interpolation, characterized by a novel pose-control model, PoseFuse3D. It effectively extracts control signals from 3D SMPL-X and fuses 2D signals, allowing precise and informative control. (ii) For evaluation, we construct CHKI‑Video, a benchmark dataset with comprehensive human‑centric annotations, which are absent in existing interpolation benchmarks. (iii) Through extensive experiments, we demonstrate that PoseFuse3D-KI delivers state‑of‑the‑art performance on our CHKI-Video benchmark with an improvement of 1.85 dB in PSNR and a reduction of 0.0796 in LPIPS.
@@ 2 Related Work
Frame Interpolation. Traditional frame interpolation methods are primarily designed for temporally adjacent frames and rely on either direct synthesis using convolutional networks [choi2020snufilm, kalluri2023flavr], or motion representations such as dynamic kernels [Niklaus_CVPR_2017, Niklaus_ICCV_2017, peleg2019net, EDSC, ding2021cdfi, lee2020adacof] and optical flows [huang2022rife, li2023amt, Kong_2022_ifrnet, siyao2021deep, hou2023video, DAIN, guo2025generalizable, jiang2018superslomo, xu2019quadratic, zhang2023emavfi]. Recent advances [feng2024trf, yang2024vibidsampler, wang2024generative] extend this task to more challenging keyframe interpolation scenarios by leveraging the generative priors of image-to-video diffusion models [blattmann2023svd, wang2025wan, rombach2022high, ho2020denoising]. These methods combine temporal forward and reverse denoising predictions in a unified process to enable interpolation. However, such approaches struggle when faced with complex articulated motions or ambiguous transitions from human keyframes. To alleviate motion ambiguity in interpolation, FCVG [zhu2024generative] introduces 2D matched lines as control signals. However, this control signal lacks the 3D geometric context required for plausible human-centric interpolation, resulting in unrealistic body shape and appearance.
@@
Pose-Guided Human Animation. Recent advances in pose-guided human animation [zhou2024realisdance, zhu2024champ, hu2024animateanyone, peng2024controlnext, zhang2023controlnet, xu2024magicanimate, wang2024disco] harness the power of diffusion models and have achieved remarkable success in generating videos from a single reference image. These methods offer flexible and precise control by incorporating enriched conditioning signals and increasingly sophisticated control mechanisms. 2D human poses, such as OpenPose [cao2017openpose] and DWPose [yang2023dwpose], are widely used in existing works [zhu2024champ, zhou2024realisdance, peng2024controlnext, zhang2023controlnet], but they are limited in capturing fine-grained geometry and motion dynamics. To address this, recent works [zhu2024champ, zhou2024realisdance] integrate 3D human parametric models [loper2023smpl, pavlakos2019smplx], which offer realistic body representation through blend shapes and skinning, resulting in more accurate and expressive human animations. Specifically, Champ [zhu2024champ] renders 3D human models into 2D proxies (e.g., normals, depth, and semantic maps) and combines them with 2D pose visualizations as control input. It operates directly on these visualizations and unifies them using a simple summation operation. To incorporate control signals into diffusion models, many approaches [wang2024disco, xu2024magicanimate, zhang2023controlnet, zhou2024realisdance] adapt ControlNet [zhang2023controlnet] for their customized control networks. Some methods [zhu2024champ, hu2024animateanyone] introduce task-specific pose guiders but often require retraining the majority of denoiser parameters. ControlNeXt [peng2024controlnext] improves control efficiency by encoding conditions with a lightweight convolutional network and injecting them via cross-normalization after the first denoising block, tuning only a minimal subset of parameters. This efficient mechanism enables robust conditioning over large-scale pre-trained video generators [wang2025wan, blattmann2023svd].
@@
In this paper, we demonstrate the advantage of combining 2D poses (DWPose [yang2023dwpose]) with 3D human models (SMPL-X [pavlakos2019smplx]) through comprehensive experiments. We present a novel pose control model that extracts unified, 3D-informed control features to provide precise guidance. Instead of relying on rendered normals and depth maps, the model directly extracts explicit 3D information from human models in 3D space, preserving richer control signals. It adopts a ControlNeXt-inspired strategy to control video diffusion models for keyframe interpolation.
@@ 3.1 PoseFuse3D
PoseFuse3D is a 3D-informed control model that provides 3D human structure and geometry guidance for plausible human interpolation. This 3D-informed guidance is injected into the base diffusion model after the first denoising block via cross-normalization [peng2024controlnext]. Internally, PoseFuse3D comprises three jointly trained components: a visual encoding module that derives control features from 3D SMPL-X [pavlakos2019smplx] renderings and 2D DWPose [yang2023dwpose] visualizations; a SMPL‑X encoder that directly embeds 3D humans and aggregates them into image conditioning maps; and a fusion module that integrates encoding streams into a unified control tensor to guide interpolation in the base VDM.
@@
Visual Encoding. This visual encoding module extracts conditional features from visualized control images, maintaining natural pixel-level alignment with the controlled video latents. To enhance the control signals, we incorporate visualizations from both DWPose and SMPL-X. While SMPL-X renderings provide rich human surface details, their keypoint information is indirect, mixed with other vertices and mesh faces. Therefore, we add DWPose visualizations to emphasize the skeletal keypoint layout for robust pose understanding. This combination was also demonstrated to be effective in prior work [zhu2024champ]. Specifically, our visual encoding module employs two parallel convolutional encoders to capture comprehensive pose information. One encoder processes DWPose visualizations to capture compact pose information, while the other handles SMPL-X renderings that retain 3D cues such as occlusion boundaries and projected shapes. Notably, to enrich semantic detail, we use SMPL-X Colored Surface [zhou2024realisdance] during rendering, which assigns a unique color to each vertex. The resulting feature maps, $E^{2D}$ and $E^{3D}$, are then passed to the fusion module for unified conditioning.
SMPL-X Encoder. Although 3D model renderings offer aligned image maps for conditioning in VDMs, the rendering operation discards parts of 3D information, particularly in occluded regions. This results in implausible interpolation for keyframes of large human motion. For example, in the first case of Figure 3, all variants relying solely on rendered controls fail to interpolate the occluded arms with correct spatial position. To enhance controllability with direct 3D information, we introduce a SMPL‑X encoder (Figure 2(b)) that processes the SMPL‑X model in 3D space and transforms it into an image conditioning feature $S^{3D}$. Specifically, a SMPL-X model [pavlakos2019smplx] is parameterized with $(\theta,\beta,\psi)$, corresponding to pose, shape, and expression. We obtain structural information by forwarding these parameters into the SMPL‑X model to generate vertex and joint coordinates $P^{3D},J^{3D}$ in 3D space and obtain their corresponding 2D coordinates $P^{2D},J^{2D}$ through projection:
@@
where $Q\in\mathbb{R}^{B\times HW\times d}$ is the flattened $d$-dimension feature of a standard 2D meshgrid extracted by a convolutional encoder. Finally, the outputs $O^{J}$ and $O^{P}$ are concatenated and passed through a downsampling block to produce the final SMPL‑X control representation $S^{3D}$, which serves as an informative and compact image embedding of the underlying 3D human structure.
Condition Fusion. The condition fusion module combines control features from both 2D and 3D signals into a unified representation to guide keyframe interpolation. For robust feature representation, we introduce a coarse-to-fine fusion strategy that progressively integrates rich geometric information from the 3D features into the compact 2D pose features. Specifically, we adopt two attention-based fusion blocks to perform this integration, where each block contains three attention layers for progressive refinement. The first layer is a self-attention module that processes the 3D features by operating on the sum of rendering encoding $E^{3D}$ and SMPL-X features $S^{3D}$. The second layer performs cross-attention, aligning the 3D features with the 2D encoding $E^{2D}$ through a spatially localized interaction scheme. Notably, we adapt the shifted window-partition strategy [liu2021swin, zhang2023emavfi] to restrict attention computation to adjacent regions, enhancing local alignment. The third layer applies temporal self-attention to capture temporal dynamic correlations in the fused representation. We use the second fusion block output as the final control signal, which is injected into the base interpolation engine to provide fine-grained, structure-aware guidance during synthesis.